# Parameter-Efficient Fine-Tuning in Large Models: A Survey of Methodologies

### v2

@@ Abstract
The large models, as predicted by scaling raw forecasts, have made groundbreaking progress in many fields, particularly in natural language generation tasks, where they have approached or even surpassed human levels. However, the unprecedented scale of their parameters brings significant computational and storage costs. These large models require substantial computational resources and GPU memory to operate. When adapting large models to specific downstream tasks, their massive parameter scale poses a significant challenge in fine-tuning on hardware platforms with limited computational power and GPU memory. To address this issue, Parameter-Efficient Fine-Tuning (PEFT) offers a practical solution by efficiently adjusting the parameters of large pre-trained models to suit various downstream tasks. Specifically, PEFT adjusts the parameters of pre-trained large models to adapt to specific tasks or domains, minimizing the introduction of additional parameters and the computational resources required. This review mainly introduces the preliminary knowledge of PEFT, the core ideas and principles of various PEFT algorithms, the applications of PEFT, and potential future research directions. By reading this review, we believe that interested parties can quickly grasp the PEFT methodology, thereby accelerating its development and innovation.
@@ I Introduction
In recent years, large pre-trained models, commonly referred to as ”large models”, have emerged as a significant advancement in the field of artificial intelligence. Due to their outstanding performance and versatility in various application contexts, these models have attracted plenty of attention and provoked much discussion. These models have impressive computing capabilities and extensive data resources, allowing them to excel in tackling intricate jobs. Within the field of natural language processing (NLP), notable interest is given to Large Language Models (LLMs). These models demonstrate remarkable ingenuity in text generation [1][2], machine translation [3][4], personalized chatbots [5][6][7], text summarization [8], sentiment analysis [9], and question-answering systems [10].
Nevertheless, the development of large models faces significant challenges and controversies. These models require substantial computational resources and data support, which can potentially jeopardize the environment and compromise privacy protection [11]. Despite their impressive performance in specific tasks, these models still have limitations and error rates that need continuous optimization and improvement [12][13][14]. When directly using large models for specific tasks, their performance often falls below desired levels. Consequently, fine-tuning large models has become a crucial method for enhancing model performance.
PEFT is a transfer learning method specifically developed to adapt the parameters of the large pre-trained models to suit new tasks and scenarios. This approach involves dynamically adjusting the model to enhance its effectiveness in performing certain tasks, taking into account the distinct features and requirements of the target task. The fine-tuning process typically entails improving the model architecture [15], optimizing parameters [16][17], and adapting learning strategies [18], among other considerations, to achieve better performance in new tasks. As the field of deep learning continues to evolve, techniques for optimizing and fine-tuning large models have also made significant advancements. Notable PEFT approaches include LoRA [19], adapter tuning [20], prefix-tuning [16], prompt-tuning [17], P-tuning [21], BitFit [22] and others. However, despite the significant achievements of large model fine-tuning techniques across several fields, there are always challenges and difficulties that need to be resolved. Overfitting mitigation, optimizing fine-tuning efficiency, and striking a learning balance between pre-training and fine-tuning tasks are a few examples of issues that need more investigation.
In recent years, hundreds of articles on PEFT have been published, with some studies offering informative overviews of the most prevalent approaches. Below is a comparative analysis of these studies.
@@
Additive Selective Reparam. Hybrid Unified NLP Vision Multi-task Diffusion Model
@@
TABLE I: Comparison with existing surveys
Ding, Ning, et al. [23] introduce a theoretical abstraction for Delta Tuning, which is analyzed from the viewpoints of optimization and optimum control. This abstraction offers a unified approach to describe the current parameter-efficient fine-tuning methods which provides a distinct perspective for future investigations. Nonetheless, while the study predominantly concentrates on NLP applications, the generalizability and efficacy of these methods in diverse domains merit additional investigation. Lialin, V, et al. [24] provide a comprehensive analysis and classification that covers a broad range of methods and compares approximately 30 approaches across five dimensions: storage efficiency, memory efficiency, computational efficiency, accuracy, and inference overhead. However, while the article primarily focuses on detailed methods with practical efficiency for fine-tuning multibillion-scale language models, the exploration of real-world application scenarios is relatively limited. Xu, Lingling, et al. [25] provide a thorough evaluation and analysis of current PEFT approaches, assessing their performance, parameter efficiency, and memory utilization within a range of NLP tasks. Nonetheless, the paper does not fully expound on the practical applications of these methodologies in actual operational environments, nor does it deeply investigate their adaptability and the domain-specific challenges they might encounter. Xin, Yi, et al. [26] offer a comprehensive overview and future directions for visual PEFT, with a systematic review of the latest advancements. While the article spans multiple visual tasks, the experiments are primarily focused on several common tasks and do not fully encompass the broader range of potential application scenarios. Han, Zeyu, et al. [27] provide a detailed classification of PEFT approaches and explores the application of PEFT techniques across various model architectures and downstream tasks, as well as the systematic design challenges of parameter-efficient fine-tuning methods. It offers researchers and engineers a comprehensive overview of PEFT approaches, but there is still room for improvement in terms of practical application coverage.
@@
  • With existing review papers, we offer a thorough and detailed overview of the fundamental knowledge related to large models and the general fine-tuning methodology. This part encompasses not only the fundamental principles, structures, and fundamental technologies of large models but also provides an in-depth overview of their practical applications in many tasks, such as natural language processing, multimodal activities, and other domains.
  • Our survey encompasses a more up-to-date review of research methodologies, highlighting the latest advancements in the field of large models. This ensures that our review is comprehensive and thorough. The scope of our review is extensive, covering multiple scenarios such as natural language processing, multimodal tasks, and vision. This provides readers with a full understanding of the current state and future prospects of large model technology.
  • After reviewing and analyzing current approaches, we suggest some innovative and future-oriented directions for further research. These areas take into account the growth potential of advanced model technology and also incorporate industry requirements and obstacles in real-world applications, proposing viable and inventive research avenues.
@@ II Preliminary
II-A Large Language Models
II-A1 Background
LLMs refer to neural language models with a large number of parameters, typically over billions of parameters. These models are built on the transformer architecture [28] and are pre-trained on vast text corpora [29]. Prior to the emergence of LLMs, the advent of transformers revolutionized the development approach for neural language models, shifting from end-to-end training to a pre-train then fine-tune paradigm. Under the pre-train fine-tune paradigm, pre-trained models can be repeatedly utilized, significantly enhancing the scalability of neural language models. Consequently, the scale of parameters is continuously growing larger. For instance, OpenAI’s GPT-1 possessed 120 million parameters, while GPT-2 boasted 1.5 billion parameters. This number surged to 175 billion for GPT-3 and soared to 1.76 trillion for the latest GPT-4 [30]. With the evident increase in neural network parameter scales, the concept of large language models has been developed.
II-A2 Emergent abilities
Research suggests that the rapid expansion of the parameter scale may lead to emergent abilities [31], which are formally defined as abilities that are not present in small models but arise in large models, constituting one of the most prominent characteristics distinguishing LLM from previous PLM. To conclude, emerging abilities can be categorized into threefolds.
@@ II-A2 Emergent abilities
II-A3 Scaling Laws of LLMs
Thanks to the exceptional scalability of the transformer architecture [28], language models also exhibit high scalability. Scaling laws for LLMs describe how the model grows and performs as the volume of training data increases.
In general, a scaling law includes four parameters, which also characterize a language model: denoted as: (1) Parameters count $N$. The number of parameters of an LLM is often associated with the number of transformer layers and the hidden size, except for some MoE LLMs. (2) Data size $D$. In LLM, this refers to the number of tokens for training. (3) Computation cost $C$. This is typically measured in terms of time and computational resources. (4) Loss $L$. The performance of training is usually evaluated by the training loss. There are two representative scaling laws for transformer LLMs.
@@ II-A3 Scaling Laws of LLMs
where $E$ denotes the loss of an ideal generative process on the test data. Furthermore, claimed by the research, the constants in this formula are $\alpha=0.34,\beta=0.28,A=406.4,B=410.7,L_{0}=1.69$. Moreover, there is a general constraint that model the relationship between $C$ and $(N,D)$: $C=6ND$, which means that it costs six FLOPs per parameter to train one token. Then it can be computed that an optimal model size and data size selection can be denoted as:
@@
II-B Prevalent LLMs
II-B1 The GPT Family
@@ II-B2 The LLaMA Family
LLaMA stands as a series of open-source LLMs developed by Meta. To date, the official release includes three versions: LLaMA, LLaMA-2, and LLaMA-3, spanning parameter scales from 3 billion to 70 billion. Beyond the weights provided by Meta, the qualities of these LLMs are further extended through supervised fine-tuning and parameter-efficient fine-tuning.
@@
LLaMA-1 [43] was released in February 2023. Although LLaMA is open-sourced and possesses fewer parameters, LLaMA-13B demonstrates significant improvements over GPT-3 (175 billion parameters) across various benchmarks. As a consequence, LLaMA has emerged as a widely adopted and exemplary base model for large language model research. LLaMA-2 was developed in in partnership with Microsoft and released half a year later. The model maintains the same architecture as the LLaMA-1 but is trained with 40% more data. LLaMA-3 was released by Meta in April 2024, offering two parameter sizes: 8B and 70B. These models underwent pre-training on approximately 15 trillion tokens of text sourced from publicly available data and are fine-tuned over 10 million human-annotated examples.
II-B3 The OpenAI o1 Family
In September 2024, OpenAI officially released a new series of large language model, o1, which trained with reinforcement learning to perform complex reasoning. o1 excels in reasoning capabilities, able to generate long chains of hidden thought processes before providing final answers, demonstrating reasoning abilities similar to humans. The release includes two versions: o1-preview and o1-mini. The o1-preview is an early iteration of the full model, while the o1-mini is a lightweight version optimized for size and speed.
II-B4 Other Representative LLMs
@@
The PaLM [46] (Pathways Language Models) is developed by Google as a collection of decoder-only LLMs. The first PaLM model was trained on a high-quality text corpus of 780 billion tokens, boasting a remarkable 540 billion parameters. Unlike prevalent LLMs which primarily utilize GPUs for training, PaLM is pre-trained with the Pathways system on 6144 TPU v4 chips to facilitate rapid and efficient training. In the following days, U-PaLM [47], FlAN-PaLM [35] and PaLM-2 were released.
II-C Multimodal Large Language Models
II-C1 MLLM: Background
Multimodal Large Language Model (MLLM), is an extension of LLM which adopts multimodal information as input such as text, sound, video, etc, to enable multiple dimensional reasoning and text generation.
Before the emergence of MLLM, significant research efforts were dedicated to multi-modality. These efforts can generally be categorized into representative and generative paradigms. An exemplary work in the representative paradigm is CLIP [48], which serves as a foundational contribution. This process yields a visual encoder  [49][50] and a text encoder, effectively establishing a bridge for downstream multimodal tasks. In contrast, generative frameworks [51][52] approach multimodal tasks by transforming them into sequence-to-sequence tasks. MLLM distinguishes itself from previous multimodal research in two key aspects. (1) Composition: MLLM is comprised of at least one LLM with billion-scale parameters. (2) Training techniques: MLLM introduces and incorporates novel training techniques derived from LLM to enhance multimodal performance.
Figure 1: Architecture of MLLM
II-C2 MLLM: Architecture

### v3

@@ Abstract
The large language models, as predicted by scaling law forecasts, have made groundbreaking progress in many fields, particularly in natural language generation tasks, where they have approached or even surpassed human levels. However, the unprecedented scale of their parameters brings significant computational and storage costs. These large language models require substantial computational resources and GPU memory to operate. When adapting large language models to specific downstream tasks, their massive parameter scale poses a significant challenge in fine-tuning on hardware platforms with limited computational power and GPU memory. To address this issue, Parameter-Efficient Fine-Tuning (PEFT) offers a practical solution by efficiently adjusting the parameters of large pre-trained models to suit various downstream tasks. Specifically, PEFT adjusts the parameters of pre-trained large language models to adapt to specific tasks or domains, minimizing the introduction of additional parameters and the computational resources required. This review mainly introduces the preliminary knowledge of PEFT, the core ideas and principles of various PEFT algorithms, the applications of PEFT, and potential future research directions. By reading this review, we believe that interested parties can quickly grasp the PEFT methodology, thereby accelerating its development and innovation.
@@ I Introduction
In recent years, large pre-trained models, commonly referred to as “large language models”, have emerged as a significant advancement in the field of artificial intelligence. Due to their outstanding performance and versatility in various application contexts, these models have attracted plenty of attention and provoked much discussion. These models have impressive computing capabilities and extensive data resources, allowing them to excel in tackling intricate jobs. Within the field of natural language processing (NLP), notable interest is given to Large Language Models (LLMs). These models demonstrate remarkable ingenuity in text generation [1, 2], machine translation [3, 4], personalized chatbots [5, 6, 7], text summarization [8], sentiment analysis [9], and question-answering systems [10].
Nevertheless, the development of large language models faces significant challenges and controversies. These models require substantial computational resources and data support, which can potentially jeopardize the environment and compromise privacy protection [11]. Despite their impressive performance in specific tasks, these models still have limitations and error rates that need continuous optimization and improvement [12, 13, 14]. When directly using large language models for specific tasks, their performance often falls below desired levels. Consequently, fine-tuning large language models has become a crucial method for enhancing model performance.
Parameter-Efficient Fine-Tuning (PEFT) is a transfer learning method specifically developed to adapt the parameters of the large pre-trained models to suit new tasks and scenarios. This approach involves dynamically adjusting the model to enhance its effectiveness in performing certain tasks, taking into account the distinct features and requirements of the target task. The fine-tuning process typically entails improving the model architecture [15], optimizing parameters [16, 17], and adapting learning strategies [18], among other considerations, to achieve better performance in new tasks. As the field of deep learning continues to evolve, techniques for optimizing and fine-tuning large language models have also made significant advancements. Notable PEFT approaches include LoRA [19], adapter tuning [20], prefix tuning [16], prompt tuning [17], P tuning [21], BitFit [22], and others. However, despite the significant achievements of large model fine-tuning techniques across several fields, there are always challenges and difficulties that need to be resolved. Overfitting mitigation, optimizing fine-tuning efficiency, and striking a learning balance between pre-training and fine-tuning tasks are a few examples of issues that need more investigation.
In recent years, hundreds of articles on PEFT have been published, with some studies offering informative overviews of the most prevalent approaches. A comparative analysis of these surveys in terms of taxonomy and application is shown in Table I.
@@
Add. Sel. Rep. Hybrid Unified NLP Vision Multi. Diffusion
@@
TABLE I: A Comparative Analysis of Survey Methodologies: Taxonomy and Application Domains, with Abbreviations: Additive (Add.), Selective (Sel.), Reparameterized (Rep.), Multi-task (Multi.) and Diffusion Model (Diffusion)
[23] introduce a theoretical abstraction for Delta Tuning, which is analyzed from the viewpoints of optimization and optimum control. This abstraction offers a unified approach to describe the current parameter-efficient fine-tuning methods which provides a distinct perspective for future investigations. Nonetheless, while the study predominantly concentrates on NLP applications, the generalizability and efficacy of these methods in diverse domains merit additional investigation. [24] provide a comprehensive analysis and classification that covers a broad range of methods and compares approximately 30 approaches across five dimensions: storage efficiency, memory efficiency, computational efficiency, accuracy, and inference overhead. However, while the article primarily focuses on detailed methods with practical efficiency for fine-tuning multibillion-scale language models, the exploration of real-world application scenarios is relatively limited. [25] provide a thorough evaluation and analysis of current PEFT approaches, assessing their performance, parameter efficiency, and memory utilization within a range of NLP tasks. Nonetheless, the paper does not fully expound on the practical applications of these methodologies in actual operational environments, nor does it deeply investigate their adaptability and the domain-specific challenges they might encounter. [26] offer a comprehensive overview and future directions for visual PEFT, with a systematic review of the latest advancements. While the article spans multiple visual tasks, the experiments are primarily focused on several common tasks and do not fully encompass the broader range of potential application scenarios. [27] provide a detailed classification of PEFT approaches and explores the application of PEFT techniques across various model architectures and downstream tasks, as well as the systematic design challenges of parameter-efficient fine-tuning methods. It offers researchers and engineers a comprehensive overview of PEFT approaches, but there is still room for improvement in terms of practical application coverage.
@@
  • This survey comprehensively reviews the latest literature PEFT, covering cutting-edge methods and related research. It establishes a theoretical framework and offers a solid knowledge base for future research.
  • We make extensive use of intuitive schematic diagrams and structured tables to elaborate on PEFT methodologies. By means of visualization, we demonstrate the complex principles of these methods, carry out comparative analyses of different approaches, and organically combine intuitiveness with systematicness, which significantly enhances the readability and academic value of the research content.
  • Breaking traditional boundaries, this survey explores PEFT in natural language processing, computer vision, multimodal fusion, and diffusion models. It uncovers application potential, offers practical guidelines, and broadens the application scope of fine-tuning technology.
@@ II Preliminary
II-A Large Language Models: Foundations and Variants
II-A1 Large Language Models
Background
LLMs refer to neural language models with a large number of parameters, typically over billions of parameters. These models are built on the transformer architecture [28] and are pre-trained on vast text corpora [29]. Prior to the emergence of LLMs, the advent of transformers revolutionized the development approach for neural language models, shifting from end-to-end training to a pre-train then fine-tune paradigm. Under the pre-train fine-tune paradigm, pre-trained models can be repeatedly utilized, significantly enhancing the scalability of neural language models. Consequently, the scale of parameters is continuously growing larger. For instance, OpenAI’s GPT-1 possessed 120 million parameters, while GPT-2 boasted 1.5 billion parameters. This number surged to 175 billion for GPT-3 and soared to 1.76 trillion for the latest GPT-4 [30].
Emergent abilities
Research suggests that the rapid expansion of the parameter scale may lead to emergent abilities [31], which are formally defined as abilities that are not present in small models but arise in large language models, constituting one of the most prominent characteristics distinguishing LLM from previous PLM. In conclusion, emerging abilities can be categorized into threefolds.
@@ II-A1 Large Language Models
Emergent abilities in large language models (LLMs) have significantly boosted various real-world applications, across fields such as natural language [36, 37, 38], healthcare [39, 40], legal [41], financial [42] and multiple scientific disciplines [43, 44]. Despite the promising emergent capabilities, there are three main limitations that restrict the further and deeper applications of LLMs. Firstly, the inconsistency across models and tasks. LLMs trained on different architectures or datasets may demonstrate emergent behavior to varying degrees. Some models might excel in certain tasks while failing to exhibit the same level of ability in others, resulting in unpredictable performance when applied to diverse real-world scenarios [45]. Secondly, the hallucinations and factual errors. LLMs often generate text that is fluent and coherent. However, they can also produce hallucinations, outputs that seem plausible but contain factual inaccuracies or misleading information [46, 47]. This tendency is particularly problematic in contexts where precise and reliable information is crucial, such as legal, medical, or scientific applications. Finally, the deficiency in deep understanding. The performance of LLMs largely stems from recognizing statistical patterns in vast datasets rather than a genuine semantic understanding of the content [46].This superficial grasp of language limits their effectiveness in tasks requiring in-depth logical reasoning and nuanced comprehension across models and tasks.
In conclusion, emergent abilities grant LLMs remarkable problem-solving capabilities, though they remain imperfect. To bridge the gap between LLMs and real-world applications, integrating traditional algorithms, expert systems, or hybrid models may be necessary to enhance reliability, accuracy, and domain-specific expertise.
Scaling Laws of LLMs
Thanks to the exceptional scalability of the transformer architecture [28], language models also exhibit high scalability. The scaling laws for LLM describe how the model grows and performs as the volume of training data increases.
In general, a scaling law includes four parameters, which also characterize a language model: (1) Parameters count $N$. The number of parameters of an LLM is often associated with the number of transformer layers and the hidden size, except for some MoE LLMs. (2) Data size $D$. In LLM, this refers to the number of tokens for training. (3) Computation cost $C$. This is typically measured in terms of time and computational resources. (4) Loss $L$. The performance of training is usually evaluated by the training loss. There are two representative scaling laws for transformer LLMs.
@@
where $E$ denotes the loss of an ideal generative process on the test data. Furthermore, claimed by the research, the constants in this formula are $\alpha=0.34,\beta=0.28,A=406.4,B=410.7,L_{0}=1.69$. Moreover, there is a general constraint that model the relationship between $C$ and $(N,D)$: $C=6ND$, which means that it costs six FLOPs per parameter to train one token. Thus, the optimal selection of model size and data size can be determined and expressed as:
@@
From the equations, scaling laws can guide decisions regarding model size. Given a fixed compute budget (e.g., 100K GPU hours), they enable predictions on whether a smaller model trained for a longer duration or a larger model trained for a shorter time would yield better performance. Additionally, scaling laws provide insight into the benefits of continued training. The diminishing returns they imply suggest that beyond a certain point, increasing compute resources may not lead to a substantial enough performance gain to justify the additional cost.
In addition, based on the statistical modeling illustrated by equation 4, one approximate estimation for Chinchilla efficient model size and training dataset size can be denoted as:
$N_{opt}$ $=0.1\,C^{0.5}$ (8)
$D_{opt}$ $=1.7\,C^{0.5}\enskip.$ (9)
This suggests that the model size and training data volume should be scaled in accordance with the available computational budget. The expected ratio of training tokens to model parameters is approximately 17:1. However, in real-world applications, this ratio is often slightly higher, as additional training data beyond the 17× scaling rule can still contribute to performance improvements when sufficient computational resources are available. For instance, GPT-2 was trained on 40B tokens with 1.5B parameters, LLaMA was trained on 1.4T tokens with 65B parameters, and DeepSeek-V3 was trained on 14.8T tokens with 0.671T parameters. While all these ratios exceed 17, they remain close to this scaling guideline.
PEFT and Sustainability of AI Research Training large models from scratch is highly energy-intensive. For example, training LLaMA-3.1 405B can demand 40 million GPU hours on H100, resulting in a substantial carbon footprint. While fully Supervised Fine-Tuning (SFT) can enhance an existing LLM using a relatively smaller set of training samples, it still requires updating the entire parameter network. In contrast, Parameter-Efficient Fine-Tuning (PEFT) methods—such as adapters or low-rank adaptations—enable fine-tuning a large pre-trained model for specific tasks by updating only a small subset of parameters (typically just 1–2% of the total). As a result, PEFT significantly reduces computational costs; for instance, a full SFT process that requires 4 million GPU hours can be reduced to 400K GPU hours or less with PEFT.
By lowering GPU usage, PEFT not only decreases energy consumption but also mitigates the environmental impact. Moreover, this reduction in compute requirements is crucial for sustainable AI research, as PEFT provides a cost-effective and efficient approach for the AI community and researchers to conduct experiments and develop new models.
II-A2 Prevalent LLMs
The GPT Family
@@ II-A2 Prevalent LLMs
LLaMA stands as a series of open-source LLMs developed by Meta. To date, the official release includes: LLaMA, LLaMA-2, and LLaMA-3.x, spanning parameter scales from 1 billion to 405 billion. Beyond the weights provided by Meta, the qualities of these LLMs are further extended through supervised fine-tuning and parameter-efficient fine-tuning.
@@
LLaMA-1 [55] was released in February 2023. Although LLaMA is open-sourced and possesses fewer parameters, LLaMA-13B demonstrates significant improvements over GPT-3 (175 billion parameters) across various benchmarks. As a consequence, LLaMA has emerged as a widely adopted and exemplary base model for large language model research. LLaMA-2 [55] was developed in partnership with Microsoft and released half a year later. The model maintains the same architecture as the LLaMA-1 but is trained with 40% more data. LLaMA-3 was released by Meta in April 2024, offering two parameter sizes: 8B and 70B. These models underwent pre-training on approximately 15 trillion tokens of text sourced from publicly available data and are fine-tuned over 10 million human-annotated examples. Subsequently, Meta released LLaMA-3.1 [56], a 405B open-sourced LLM, which focuses on improving text generation capabilities and achieves performance comparable to leading models like GPT-4. Then, in September 2024, LLaMA-3.2 was released, introducing both vision models (11B and 90B) and lightweight text-only models (1B and 3B) for mobile device use. LLaMA-3.2 marked Meta’s first open-source AI model capable of processing both images and text, broadening the scope of potential applications. The smaller models were designed for efficient performance on mobile devices, promoting wider adoption in edge computing scenarios.
The OpenAI o1 Family
In September 2024, a new series of large language model, OpenAI-o1111https://openai.com/index/introducing-openai-o1-preview/ [57], excels in complex reasoning tasks, using Chain-of-Thought (CoT) reasoning to outperform GPT-4o in areas like math, coding, and science. The release includes two versions: o1-preview and o1-mini. The o1-preview is an early iteration of the full model, while the o1-mini is a lightweight version optimized for size and speed. When solving problems, o1 uses the CoT [^2] https://openai.com/index/learning-to-reason-with-llms/ strategy like human deep thinking. Reinforcement learning helps o1 refine its thinking and strategies, find and correct errors, break down complex steps, and change approaches when necessary, improving reasoning. The reward model combines text and number scores for evaluation.
Then previewed in December 2024, OpenAI o3-mini [^3] https://openai.com/index/openai-o3-mini/, the newest, most cost-efficient model was offically released in January 2025, which provides a specialized alternative for technical domains requiring precision and speedwhich. It delivers exceptional STEM capabilities—with particular strength in science, math, and coding—all while maintaining the low cost and reduced latency of OpenAI o1-mini.
The DeepSeek Family
DeepSeek-LLM is a newly established LLM series that has garnered significant attention from both academia and industry. Developed by the company DeepSeek, the first version, DeepSeek-V1 [58], was trained on 2 trillion tokens and released in January 2024, featuring two core models: 7B and 67B, along with their respective chat variants. In the same month, DeepSeek introduced DeepSeek-MoE (Mixture of Experts) [59] 16B, which delivers performance comparable to LLaMA 2 7B while requiring only 40% of the computational cost. This model introduces an innovative Mixture of Experts (MoE) architecture, integrating shared expert isolation with fine-grained expert segmentation. Additionally, it incorporates a novel load-balancing strategy that optimizes both expert and device balance, enhancing computational efficiency. They made significant progress with DeepSeek-V2 [60], a large MoE-LLM trained on 8.1 trillion tokens, featuring 2 shared experts, 160 routed experts, and 236 billion parameters. This version introduced Multi-head Latent Attention (MLA), which significantly reduces GPU memory consumption while maintaining the same level of precision. It outperforms the widely used Grouped-Query Attention (GQA) strategy adopted by LLaMA 3. Subsequently, they released DeepSeek-V3 [38] in December 2024. Building upon V2, the V3 model introduces its Multi-Token Prediction (MTP) approach and an Auxiliary-Free Load Balancing strategy to further enhance efficiency. Additionally, it integrates DualPipe [61], cross-node all-to-all communication techniques, and a minimal-overhead memory-saving strategy, achieving a groundbreaking industrial milestone—training a 671B-parameter MoE-LLM with FP8 precision. The performance of the DeepSeek-V3 model is remarkable, achieving state-of-the-art (SOTA) results among all open-source LLMs and demonstrating performance comparable to GPT-4o and Claude 3.5 Sonnet. Moreover, it offers significant advantages in training and inference costs, requiring less than 10% of the training cost of LLaMA 3-405B and only 9% of the inference cost of Claude 3.5 Sonnet, revolutionizing the development of industrial LLMs. Then, DeepSeek released R1 [62], a reinforcement learning-focused model leveraging the Group Relative Policy Optimization (GRPO) [63] algorithm. R1 delivers performance comparable to OpenAI-o1 in mathematical and logical reasoning tasks, while requiring only 2% of the computational cost, marking a major breakthrough in efficiency and scalability.
The Claude Family
Claude [64] represents a series of conversational AI models developed by Anthropic, designed with a focus on safety, helpfulness, and natural language understanding. This family includes Claude 1, Claude 2, Claude 2.1, Claude 3 Opus, Claude 3 Sonnet, and Claude 3 Haiku.
Claude 1 marked the initial release of Anthropic’s conversational AI, introducing the concept of Constitutional AI to the field. Claude 2 and its subsequent update, Claude 2.1, brought significant improvements in language understanding, context retention, and response coherence. These versions demonstrated enhanced capabilities in handling complex queries and maintaining longer, more contextually rich conversations.
Claude 3 models (Opus, Sonnet, and Haiku) represent the latest advancements in the Claude family, each tailored for distinct applications. Opus, the most advanced model, integrates cutting-edge multimodal capabilities, enabling it to process both textual and visual inputs with deep reasoning and high-level comprehension, excelling in complex problem-solving tasks. Sonnet, optimized for efficiency and speed, is ideal for scenarios requiring rapid, precise, and contextually appropriate replies. Haiku prioritizes simplicity and elegance, delivering concise, poetic, and highly relevant responses, making it particularly well-suited for creative and literary applications. Together, these models set new benchmarks for AI-driven interaction and analytical reasoning.
Each model in the Claude family is continuously refined to improve performance, safety, and alignment with user needs, ensuring that they remain at the forefront of conversational AI technology.
The Gemini Family
Gemini [65] constitutes a series of multimodal Transformer-based language models, developed by Google DeepMind. This family includes Gemini 1, Gemini 1.5, and Gemini 2, each introducing significant advancements in multi-modal understanding, long-context reasoning, and integration with Google’s ecosystem. Unlike GPT family models, which initially focused on text generation, Gemini models were designed from the ground up to be native multimodal models, enabling seamless processing of text, images, audio, and video. Gemini 1 marked Google’s transition from its Bard chatbot to a more advanced multimodal LLM, introducing cross-modal reasoning and excelling in mathematical problem-solving, coding, and knowledge retrieval, though it faced limitations in real-world usability. Gemini 1.5 introduced a 1 million-token context window, significantly improving long-document processing, dialogue coherence, and complex multi-step reasoning. Additionally, it implemented memory capabilities, allowing it to retain user-specific context across interactions. The latest version, Gemini 2 further enhanced reasoning, tool integration, and inference speed, introducing a “Flash Thinking” mode that enables intermediate reasoning steps for improved transparency. It also deepened integration with Google Search, Docs, and other productivity tools, optimizing it for real-world applications.
Other Representative LLMs
@@
The PaLM [68] (Pathwaysutilized Language Models) is developed by Google as a collection of decoder-only LLMs. The first PaLM model was trained on a high-quality text corpus of 780 billion tokens, boasting a remarkable 540 billion parameters. Unlike prevalent LLMs which primarily utilize GPUs for training, PaLM is pre-trained with the Pathways system on 6144 TPU v4 chips to facilitate rapid and efficient training. In the following days, U-PaLM [69], FlAN-PaLM [35] and PaLM-2 were released.
II-A3 Multimodal Large Language Models
MLLM: Background
Multimodal Large Language Model (MLLM), is an extension of LLM which adopts multimodal information as input such as text, sound, video, etc. to enable multiple dimensional reasoning and text generation.
Before the emergence of MLLM, significant research efforts were dedicated to multi-modality. These efforts can generally be categorized into representative and generative paradigms. An exemplary work in the representative paradigm is CLIP [70], which serves as a foundational contribution.
This process yields a visual encoder  [71][72] and a text encoder, effectively establishing a bridge for downstream multimodal tasks. In contrast, generative frameworks [73][74] approach multimodal tasks by transforming them into sequence-to-sequence tasks. MLLM distinguishes itself from previous multimodal research in two key aspects. (1) Composition: MLLM is comprised of at least one LLM with billion-scale parameters. (2) Training techniques: MLLM introduces and incorporates novel training techniques derived from LLM to enhance multimodal performance.
Figure 1: Architecture of MLLM: This figure shows a common architecture and workflow of an MLLM.
MLLM: Architecture